OpenStudy (anonymous):

What is the rule for the sequence 324, 108, 36, 12, ? Subtract 12 from the previous number Divide the previous number by 2 Subtract 216 from the previous number Divide the previous number by

OpenStudy (anonymous):

check!!

OpenStudy (anonymous):

is \(324-12=108\) NO

OpenStudy (anonymous):

is \(324\div 2=108\) ? NO

jagr2713 (jagr2713):

ohoho you multiply by 3 :D

OpenStudy (anonymous):

the correct answer is the one that is missing

jagr2713 (jagr2713):

12*3=36 36*3=108 108*3=324

jagr2713 (jagr2713):

wo0w its missing in the answer choices ??/

OpenStudy (mathstudent55):

Try the first two numbers of the sequence with each choice. If it doesn't work, that choice is not it, and move to the next choice. If it works, try the next pair of numbers on the choice.

OpenStudy (mathstudent55):

@jagr2713 Do you read form right to left or from left to right?

jagr2713 (jagr2713):

The nth term is 4(-3)ⁿ⁻¹ 1st term: 4(-3)⁰ = 4 2nd term: 4(-3)¹ = -12 3rd term: 4(-3)² = 36 etc. You want the sixth term.

jagr2713 (jagr2713):

i did it from 12 to 36

jagr2713 (jagr2713):

but i think it is The nth term is 4(-3)ⁿ⁻¹ 1st term: 4(-3)⁰ = 4 2nd term: 4(-3)¹ = -12 3rd term: 4(-3)² = 36 etc. You want the sixth term.

OpenStudy (anonymous):

so the awnswer is?.........................

OpenStudy (mathstudent55):

What problem are you solving?

jagr2713 (jagr2713):

like i dont even know right now lol. my brain is fried

OpenStudy (mathstudent55):

@sjman1 Choice A: Subtract 12 from the previous number 324 - 12 = 312 Is 312 = 108 ?

OpenStudy (anonymous):

Which description is true about the expression 7 plus x plus 4 times w minus 9? 5 terms 1 variable term 1 coefficient 2 constant terms

OpenStudy (mathstudent55):

Are you done with the first problem?

OpenStudy (mathstudent55):

7 + x + 4w - 9

OpenStudy (mathstudent55):

Terms are separated by plus or minus. Variables are the letters. Coefficients are the numbers that multiply variables or are alone when there is no variable. Constant term is a number alone without variable.